OpenStudy (anonymous):

Solve –3(2x – 4) = 24 x = -2 right?

jimthompson5910 (jim_thompson5910):

let's check that –3(2x – 4) = 24 –3(2(-2) – 4) = 24 -3(-4-4) = 24 -3(-8) = 24 24 = 24 So you are correct